### v1

@@ 1 Introduction
To secure system networks and privacy in deployment against cyber risks, an intrusion detection system (IDS) by analyzing and monitoring network traffic is one of the most efficient solutions (Sharma et al., 2023; Idrissi et al., 2023), these approaches are also known as anomaly detection systems in some research. Modern IDSs that leverage machine learning (ML) techniques to detect unseen threats in IoT networks have been studied and adopted widely (Sharma et al., 2023; Vu et al., 2022; Wang et al., 2023). In addition to developing effective data learning models, it is also very important to choose and define learning and deployment strategies to face new challenges when IoT networks continuously scale up. Traditional ML methods work in a centralized paradigm, in which one model is trained on a single server using all data collected from connected agents (Guendouzi et al., 2023). In the modern IoT context, this approach depicts some challenges such as computational resource requirements for training and serving models; data latency when the data transmission distance may be hundreds, even thousands of miles from the server; data transfer cost; particularly data privacy preservation issues due to data leakage when sharing data among agents and server (Abdulrahman et al., 2021). Therefore, Federated Learning (FL) has emerged recently as a cutting-edge approach to solving these problems. Rather than collecting data, training, and serving ML models in a single high-performance computer, FL takes advantage of the computing capabilities of the network client in the distributed mechanism to train the local model itself and protect the local data from being accessed directly from outside (Abdulrahman et al., 2021). Then, the local models will be aggregated to a global model in the server for serving by some methodologies.
@@
To tackle these limitations, in this article, we aim to propose FedMSE, a collaborative IoT network intrusion detection approach leveraging the Autoencoder-based feature representation method, Mean Squared Error, and FL. The source code was published on Github  [^1] https://github.com/dino-chiio/fedmse together with the dataset prepared for this paper. FedMSE enables the strength of identifying unknown cyber threats of semi-supervised learning techniques, and the secure learning process of FL. Each IoT gateway independently trains an intrusion detection model on its locally processed network data without sharing the data with a central server. We use a hybrid approach, constructed by Shrink Autoencoder (SAE) as a feature representation method and Centroid (CEN) as a one-class classifier (Cao et al., 2019), to improve the performance of intrusion detection in local gateways. Then, we develop a novel federated aggregation algorithm based on Mean Squared Error to enhance the accuracy, efficiency, and robustness of decentralized learning. This new method considers the contribution of local models in the training process and prioritizes more accurate models. The experimental evaluation demonstrates the enhancement in IoT network intrusion detection of our proposed approach.
@@ 2.2 Autoencoder
In many applications, especially in network intrusion detection, AEs can be used for feature dimension reduction by compressing the original data into a lower-dimension space. The most important characteristics of original complex data can be extracted and represented more effectively in the latent layer, making other machine learning models learn data patterns more easily.
@@ 3 Related work
As mentioned before, existing network intrusion detection approaches are facing some challenges. The scarcity of labeled anomalous data is one of the major issues in advanced intrusion detection. Despite the vast amount of network traffic data generated by modern and complex IoT networks, there is a notable lack of labeled abnormal data due to privacy issues, limitations of expert knowledge and labeling expenses (Cao et al., 2019). Anomalous network traffic includes computer communication related to system weaknesses and vulnerabilities. Thus, administrators tend to keep the data private to protect their network or client’s sensitive information. Moreover, IoT networks generate a huge amount of network data and intrusion behaviors become more sophisticated over time, leading to being more challenging to collect, process, and label anomalous traffic in real-world networks that require a significant amount of time and expert knowledge. This absence makes it difficult to use supervised machine learning models to learn the characteristics of available intrusive data. To address this, some semi-supervised learning techniques, such as OCSVM (Zhang et al., 2015), and LOF (Breunig et al., 2000), can be used to model normal data. Especially, AEs have been applied widely in network intrusion detection due to the power of modeling complex characteristics of network data (Cao et al., 2019; Meidan et al., 2018; Idrissi et al., 2023). By modeling the normal data, new data points will be classified based on the reconstruction error of these ones, a higher error suggesting a higher probability of being an anomaly.
@@
One of the fundamental challenges in FL is the handling of non-IID (non-Independent and Identically Distributed) data (McMahan et al., 2017). In traditional centralized machine learning, data is typically assumed to be IID, which means that each data point is independent of the others and is drawn from the same distribution. In FL, however, the data on each client device is inherently linked to the user’s behavior and usage patterns. Some devices will generate large amounts of data while others contribute very little, resulting in heterogeneity in IoT networks (Sattler et al., 2020). An IoT network can be divided into multiple subnetworks, each managed by an IoT gateway, also known as an edge server, and located in a different area. The training data size and distribution of each subnetwork typically vary significantly due to differences in network traffic generated by their specific local environments and applications. This heterogeneity can lead to local models that are poorly representative of the overall population, making the global model work ineffectively in intrusion detection. Addressing this issue requires careful consideration of how to weigh contributions from different gateways to ensure that the global model remains fair and representative of all gateways. Both FedAvg and FedProx, mentioned in Section 2, update the global model based on the size of the training data held by local clients. This makes them face challenges in complex distributed networks with dynamic environments where participant availability and data characteristics can vary significantly. In this case, the client that has a large number of data samples may not represent strongly for the whole network.
@@
In federated IoT applications, it is also crucial to consider some resource constraints aspects such as communication overhead, and computational expense (McMahan et al., 2017; Sattler et al., 2020). Distributed training requires a massive number of clients to participate in the training process, increasing the communication throughput and time consumption. IoT devices have small computational capacity and available resources for training and predicting simultaneously are limited. More accurate approaches require more complex computational activities, leading to an increase in resource consumption and incident response time. Therefore, choosing a lightweight approach and robust learning strategy is essential in federated IoT intrusion detection.
@@ 4.2 Mean Squared Error-based model aggregation
During training a machine learning model, it is necessary to control the convergence of the model on both local and global sides to prevent under-fitting and over-fitting problems. In the data preprocessing phase, the server will collect a small processed dataset from all gateways in the network, called development data, that includes normal network data only to validate the global model convergence in each training round. This dataset contains data of all gateways with a uniform distribution. That means that all gateways have the same amount of normal data in this set and the representativeness of the whole IoT network will be ensured. Using processed data also keeps the sensitive information private in local gateways.
@@
Figure 2 and Algorithm 2 show the MSEAvg principle and pseudo-code for the whole operation. When all participants have completed the local training process and sent the optimal weights to the server, each local model uses the development data as input, returns the reconstruction error calculated by the mean squared error of input and output, and is assigned a weight in the updating process based on this error. The smaller error implies a better model for learning normal data, which means a higher weight. After that, the global model will be aggregated by using the Equation 5:
$W_{global}=\frac{\sum_{i=1}^{n}1/\alpha_{i}\cdot W_{i}}{\sum_{i=1}^{n}1/\alpha_{i}}$ (5)
where $\alpha_{i}$ is the update weight based on MSE loss; $W_{i}$ is the local model weights.
@@ 5.1.1 IoT network construction
To mimic the IoT network for this evaluation, a small partition of the N-BaIoT dataset, with only normal data, is used to make experiments close to practical scenarios. N-BaIoT dataset (Meidan et al., 2018) is designed to aid in the detection of IoT botnet attacks. The dataset was collected from nine commercial IoT devices that were intentionally infected with two prevalent IoT-based botnets: Mirai and Gatfyt. The details of this dataset are described in Table 1.
Table 1: N-BaIoT network traffic dataset. Normal, Gatfyt and Mirai columns show the number of data samples for normal, Gatfyt botnet and Mirai botnet behaviour in each IoT device, respectively.
ID Device name Normal Gatfyt Mirai
@@ 8 Ennio_Doorbell 39100 316400
Figure 4 shows the experimental data allocation. We consider a 10-gateway IoT network, which means that there are 10 IoT gateways in the network. The gateway data is selected from N-BaIoT data using random algorithms and constraints to ensure the non-IID characteristics of training data. We use the Dirichlet distribution $Dir_{n}(\alpha)$ with $n$ is the number of gateways and $\alpha$ as the concentration parameter and Jensen-Shannon measurement (JS) to control the non-IIDness of networks. To simulate a homogeneous setting, for each device type $k$, we sample $p_{k}\sim Dir_{10}(1000)$ and allocate the proportion data $p_{k,j}$ of device $k$ for the gateway $j$, and the measurement is $\text{{JS}}=0.01$. We do the same process for the heterogeneous setting with the concentration parameter $\alpha=0.1995$ and $\text{{JS}}=0.83$. In the testing phase, some new devices will be added to the network to simulate the dynamics of IoT networks.
@@ 5.1.2 Hyperparameters setting
This research studies the accuracy of Autoencoder and SAE-CEN models under FedAvg, FedProx, and MSEAvg federated learning algorithms. In intrusion detection, it is challenging to determine a threshold that accurately classifies a data point as either normal or anomalous. Thus, the Area Under the ROC Curve (AUC) metric (Huang and Ling, 2005) is adopted to evaluate these models at different thresholds. We ran each experiment five times and calculated the mean accuracy and standard deviation of the AUC values.
Due to the lack of anomalies during training, the experiment hyperparameters can not be tuned, which is one of the considerable challenges of this work. We set up them using common values. The mini-batch size value is chosen as $12$, learning rate $lr=0.00001$, local epoch $I=100$, number of global round $E=20$, $\mu=0.001$ for the FedProx proximal term. In each round, we select half of the gateways to join the training process based on the majority rule. For the SAE shrink parameter, we set the value of 10 as Cao et al. (2019) for balancing shrink loss and MSE loss and employ the Adam optimizer (Kingma and Ba, 2014) along with early stopping techniques (Prechelt, 2012) to train these local models. We split the local training data into 40% for training, 10% for validation, 40% for the development dataset, and 10% for adding to testing set. The early stopping techniques are also used in the global server to control the convergence of the global model on the development dataset. For the number of neurons in the latent layer of the Autoencoder-based model, we configure based on rules of thumb as mentioned in Cao et al. (2016) with the value $m=\left[1+\sqrt{n}\right]$, where $n$ is the original space dimension.
@@ 5.2 Experiment 1: Federated intrusion detection performance
We do the first experiment to examine the detection ability of our approach in every network gateway. Table 2 and Figures 6, 6 present the AUCs obtained by the Autoencoder and SAE-CEN models under three aggregation algorithms. The results indicate that the larger non-IIDness setting poses greater challenges for machine learning models in detecting anomalies, leading to a decrease in accuracy and less consistency among gateways compared to the low non-IIDness scenario.
@@
Figure 6: Accuracy for federated intrusion detection in the high non-IDDness setting
@@ 5.3 Experiment 2: Effects of gateway selection ratio
Table 3 and Figure 9, 9 show the average accuracy of comparison candidates under different settings of the gateway selection ratio. The results indicate that the gateway selection ratio affects the consistency among gateways, and FedMSE still shows outstanding power. This is evidenced by changes in the standard deviation and the complexity of algorithms although the mean accuracy fluctuates slightly and all FedAvg, FedProx, and MSEAvg help SAE-CEN have comparable accuracy.
@@
5.4 Experiment 3: Accuracy for different IoT network sizes
Due to IoT networks changing in size extensively from the initial state during use time, it is necessary to choose a robust model for the dynamic characteristics of networks. This experiment is performed to investigate our approach’s robustness in different network scales. Table 4 and Figures 11 and 11 show the models’ average performance under different settings of IoT network scale with the gateway ratio of 50% and the similar hyperparameters to Experiment 1.
@@ 6 Conclusion
However, the approach has several limitations, including sensitivity to hyperparameters and the need for more efficient gateway selection strategies in federated learning settings with heterogeneous data. The MSEAvg algorithm also increases computational costs due to additional server-side computations. Future work should focus on implementing an experimental end-to-end framework to evaluate performance in practical networks, integrating techniques like hierarchy learning to enhance the aggregation algorithm, and optimizing hyperparameters for real-world applications. Addressing these challenges can further refine the proposed framework to better meet the evolving security needs of IoT networks.

### v2

@@ 1 Introduction
To secure system networks and privacy in deployment against cyber risks, an intrusion detection system (IDS) by analyzing and monitoring network traffic is one of the most efficient solutions (Sharma et al., 2023; Idrissi et al., 2023), these approaches are also known as anomaly detection systems in some research. Modern IDSs that leverage machine learning (ML) techniques to detect unseen threats in IoT networks have been studied and adopted widely (Sharma et al., 2023; Vu et al., 2022; Wang et al., 2023). They can be categorized into three main types: supervised learning, semi-supervised learning, and unsupervised learning (Chandola et al., 2009). Supervised learning works with the assumption that all available data is labeled for building the model. Any new data are fed into the trained model and classified as normal or abnormal. The semi-supervised learning methods suppose that only normal data is available to construct the model. They infer the unseen data based on the deviation from normal data. The last type is unsupervised learning, which does not require labeled training data and works under the implicit assumption that the normal data points are far from anomalous ones. Semi-supervised and unsupervised learning techniques are often used more widely in intrusion detection due to their ability to detect unknown anomalies. In addition to developing effective data learning models, it is also very important to choose and define learning and deployment strategies to face new challenges when IoT networks continuously scale up. Traditional ML methods work in a centralized paradigm, in which one model is trained on a single server using all data collected from connected agents (Guendouzi et al., 2023). In the modern IoT context, this approach depicts some challenges such as computational resource requirements for training and serving models; data latency when the data transmission distance may be hundreds, even thousands of miles from the server; data transfer cost; particularly data privacy preservation issues due to data leakage when sharing data among agents and server (Abdulrahman et al., 2021), especially in network intrusion detection. Network data packets contain extremely sensitive data information, such as device addresses, connection sessions, personal user information, system vulnerabilities, etc. The continuous exchange and centralized storage of this information in computer networks increases the risk of data leakage, leading to serious information incidents. Therefore, Federated Learning (FL) has emerged recently as a cutting-edge approach to solving these problems. Rather than collecting data, training, and serving ML models in a single high-performance computer, FL takes advantage of the computing capabilities of the network client in the distributed mechanism to train the local model itself and protect the local data from being accessed directly from outside (Abdulrahman et al., 2021). Then, the local models will be aggregated to a global model in the server for serving by some methodologies.
@@
To tackle these limitations, in this article, we aim to propose FedMSE, a collaborative IoT network intrusion detection approach leveraging the Autoencoder-based feature representation method, Mean Squared Error, and FL. The source code was published on Github  [^1] https://github.com/dino-chiio/fedmse together with the dataset prepared for this paper. FedMSE enables the strength of identifying unknown cyber threats of semi-supervised learning techniques and the secure learning process of FL. Each IoT gateway independently trains an intrusion detection model on its locally processed network data without sharing the data with a central server. We use a hybrid approach, constructed by Shrink Autoencoder (SAE) as a feature representation method and Centroid (CEN) as a one-class classifier (Cao et al., 2019), to improve the performance of intrusion detection in local gateways. Then, we develop a novel federated aggregation algorithm based on Mean Squared Error to enhance the accuracy, efficiency, and robustness of decentralized learning. This new method considers the contribution of local models in the training process and prioritizes more accurate models. The experimental evaluation demonstrates the enhancement in IoT network intrusion detection of our proposed approach.
@@ 2.2 Autoencoder
AEs are particularly useful for intrusion detection due to their ability to capture the underlying structure of normal data during training. Once trained on normal data, the AE will produce low reconstruction errors (Equation 2) for inputs similar to the training data. Nevertheless, the AE finds it difficult to correctly recreate data that is anomalous, that is, data that differs greatly from the patterns discovered during training. Reconstruction errors can be used as a metric to classify anomalies. Higher errors indicate higher probabilities for anomalous.
In many applications, especially in network intrusion detection, AEs can be used for feature dimension reduction by compressing the original data into a lower-dimension space. The most important characteristics of original complex data can be extracted and represented more effectively in the latent layer, not only making other machine learning models learn data patterns more easily but also reducing the computational complexity in detection.
@@ 3 Related work
As mentioned before, existing network intrusion detection approaches are facing some challenges. The scarcity of labeled anomalous data is one of the major issues in advanced intrusion detection. Despite the vast amount of network traffic data generated by modern and complex IoT networks, there is a notable lack of labeled abnormal data due to privacy issues, limitations of expert knowledge and labeling expenses (Cao et al., 2019). Anomalous network traffic includes computer communication related to system weaknesses and vulnerabilities. Thus, administrators tend to keep the data private to protect their network or client’s sensitive information. Moreover, IoT networks generate a huge amount of network data and intrusion behaviors become more sophisticated over time, leading to being more challenging to collect, process, and label anomalous traffic in real-world networks that require a significant amount of time and expert knowledge. This absence makes it difficult to use supervised machine learning models to learn the characteristics of available intrusive data. To address this, some semi-supervised learning techniques, such as OCSVM (Zhang et al., 2015), and LOF (Breunig et al., 2000), can be used to model normal data. Especially, AEs have been applied widely as a state-of-the-art approach in network intrusion detection due to the power of modeling complex characteristics of network data (Cao et al., 2019; Meidan et al., 2018; Idrissi et al., 2023). By modeling the normal data, new data points will be classified based on the reconstruction error of these ones, a higher error suggesting a higher probability of being an anomaly.
@@
One of the fundamental challenges in FL is the handling of non-Independent and Identically Distributed (non-IID) data (McMahan et al., 2017). In traditional centralized machine learning, data is typically assumed to be Independent and Identically Distributed (IID), which means that each data point is independent of the others and is drawn from the same distribution. In FL, however, the data on each client device is inherently linked to the user’s behavior and usage patterns. Each device operates within its unique local environment, causing its generated data to have different statistical characteristics, resulting in heterogeneity in IoT networks (Sattler et al., 2020). An IoT network can be divided into multiple subnetworks, each managed by an IoT gateway, also known as an edge server, and located in a different area. The training data size and distribution of each subnetwork typically vary significantly due to differences in network traffic generated by their specific local environments and applications. This heterogeneity can lead to local models that are poorly representative of the overall population, making the global model work ineffectively in intrusion detection. Addressing this issue requires careful consideration of how to weigh contributions from different gateways to ensure that the global model remains fair and representative of all gateways. Both FedAvg and FedProx, mentioned in Section 2, update the global model based on the size of the training data held by local clients. This makes them face challenges in complex distributed networks with dynamic environments where participant availability and data characteristics can vary significantly. In this case, the client that has a large number of data samples may not represent strongly for the whole network.
@@
In federated IoT applications, it is also crucial to consider some resource constraints aspects such as communication overhead, and computational expense (McMahan et al., 2017; Sattler et al., 2020). Distributed training requires a massive number of clients to participate in the training process, increasing the communication throughput and time consumption. IoT devices have small computational capacity and available resources for training and predicting simultaneously are limited. More accurate approaches require more complex computational activities, leading to an increase in resource consumption and incident response time. Therefore, choosing a lightweight approach and robust learning strategy is essential in federated IoT intrusion detection. In some research, AEs were used as a feature manipulation method to extract the most important patterns and compress high-dimensional data to more optimal space, such as Cao et al. (2019), Nguyen et al. (2024). This not only helps reduce the computational complexity but also increases the accuracy and the response time to incidents.
@@ 4.2 Mean Squared Error-based model aggregation
During training a machine learning model, it is necessary to control the convergence of the model on both local and global sides to prevent under-fitting and over-fitting problems. In the data preprocessing phase, the server will collect a small processed dataset from all gateways in the network, called development data, that includes normal network data only to validate the global model convergence in each training round. Each gateway randomly selects the same amount of normal data as the others, and these subsets are then combined to form the development dataset, ensuring the representativeness of the whole IoT network. Using processed data also keeps the sensitive information private in local gateways.
@@
Figure 3 and Algorithm 2 show the MSEAvg principle and pseudo-code for the whole operation. When all participants have completed the local training process and sent the optimal weights to the server, each local model uses the development data as input, returns the reconstruction error calculated by Mean Squared Error (MSE) of input and output, and is assigned a weight in the updating process based on this error. The smaller error implies a better model for learning normal data, which means a higher weight. After that, the global model will be aggregated by using the Equation 5:
$W_{global}=\frac{\sum_{i=1}^{n}\alpha_{i}\cdot W_{i}}{\sum_{i=1}^{n}\alpha_{i}}$ (5)
where $W_{i}$ is the local model weights; $\alpha_{i}$ is the update weight based on MSE loss and calculated by Equation 6:
$\alpha_{i}=\frac{1}{MSE_{i}}$ (6)
where $MSE_{i}$ is the mean squared error of local model $i$ with the development dataset.
@@
The MSEAvg algorithm enhances the global federated Autoencoder-based models by prioritizing contributions from local models with superior reconstruction accuracy, as models with lower reconstruction errors on the development dataset better represent normal network data. This approach ensures a robust and reliable global model by aligning the aggregation process with the primary objective of Autoencoder-based models in anomaly detection. In addition, MSE is more sensitive in model aggregation than alternatives such as Mean Absolute Error (MAE), which helps reduce the effect of low-performing local models. MSE squares the reconstruction errors, which amplifies the effect of larger errors during the weight calculation more than MAE. Poor-performing local models with high reconstruction errors are assigned much smaller weights. Meanwhile, well-performing local models with lower errors contribute more significantly to the global model. This helps the global model be more robust.
Unlike conventional algorithms such as FedAvg and FedProx, which weights contributions based on dataset size, MSEAvg employs a performance-based criterion. This innovation is particularly effective in handling non-IID data. By evaluating local models using a representative development dataset, MSEAvg ensures fair assessment regardless of training data heterogeneity. This performance-based weighting reduces the risk of over-fitting to data-rich gateways, thereby enhancing the generalizability and reliability of the global model.
@@ 5.1.1 IoT network construction
To mimic the IoT network for this evaluation, a small partition of the N-BaIoT dataset, with only normal data, is used to make experiments close to practical scenarios. N-BaIoT dataset (Meidan et al., 2018) is designed to aid in the detection of IoT botnet attacks. The dataset was collected from nine commercial IoT devices that were intentionally infected with two prevalent IoT-based botnets: Mirai and Gafgyt. The details of this dataset are described in Table 1.
Table 1: N-BaIoT network traffic dataset. Normal, Gafgyt and Mirai columns show the number of data samples for normal, Gafgyt botnet and Mirai botnet behaviour in each IoT device, respectively.
ID Device name Normal Gafgyt Mirai
@@ 5.1.2 Hyperparameters setting
This research studies the accuracy of Autoencoder and SAE-CEN models under FedAvg, FedProx, and MSEAvg federated learning algorithms. In intrusion detection, it is challenging to determine a threshold that accurately classifies a data point as either normal or anomalous. Thus, the Area Under the ROC Curve (AUC) metric (Huang and Ling, 2005) is adopted to evaluate these models at different thresholds. We ran each experiment five times and calculated the mean accuracy and standard deviation of the AUC values, following the research in Rey et al. (2022) that evaluated on N-BaIoT dataset.
Due to the lack of anomalies during training, the experiment hyperparameters can not be tuned, which is one of the considerable challenges of this work. We set up them using common values. The mini-batch size value is chosen as $12$, learning rate $lr=0.00001$, local epoch $I=100$, number of global round $E=20$, $\mu=0.001$ for the FedProx proximal term. In each round, we randomly select half of the gateways to join the training process based on the majority rule. For the SAE shrink parameter, we set the value of 10 as Cao et al. (2019) for balancing shrink loss and MSE loss and employ the Adam optimizer (Kingma and Ba, 2014) along with early stopping techniques (Prechelt, 2012) to train these local models. We split the local training data into 40% for training, 10% for validation, 40% for the development dataset, and 10% for adding to testing set. To ensure all gateways have an equal amount of development data, we use the size of the gateway that has the minimum number of development data samples. The early stopping techniques are also used in the global server to control the convergence of the global model on the development dataset. For the number of neurons in the latent layer of the Autoencoder-based model, we configure based on rules of thumb as mentioned in Cao et al. (2016) with the value $m=\left[1+\sqrt{n}\right]$, where $n$ is the original space dimension.
@@ 5.2 Experiment 1: Federated intrusion detection performance
Figure 7: Accuracy for federated intrusion detection in the high non-IIDness setting
We do the first experiment to examine the detection ability of our approach in every network gateway. Table 2 and Figures 6, 7 present the AUCs obtained by the Autoencoder and SAE-CEN models under three aggregation algorithms. The results indicate that the larger non-IIDness setting poses greater challenges for machine learning models in detecting anomalies, leading to a decrease in accuracy and less consistency among gateways compared to the low non-IIDness scenario.
@@ 5.3 Experiment 2: Effects of gateway selection ratio
Table 3 and Figure 9, 10 show the average accuracy of comparison candidates under different settings of the gateway selection ratio. The results indicate that the gateway selection ratio affects the consistency among gateways, and FedMSE still shows outstanding power. This is evidenced by changes in the standard deviation and the complexity of algorithms although the mean accuracy fluctuates slightly and all FedAvg, FedProx, and MSEAvg help SAE-CEN have comparable accuracy.
@@ 5.4 Experiment 3: Accuracy for different IoT network sizes
Due to IoT networks changing in size extensively from the initial state during use time, it is necessary to choose a robust model for the dynamic characteristics of networks. This experiment is performed to investigate our approach’s robustness in different network scales. Table 4 and Figures 11 and 12 show the models’ average performance under different settings of IoT network scale with the gateway ratio of 50% and the similar hyperparameters to Experiment 1.
@@ 6 Conclusion
However, the approach has several limitations, including sensitivity to hyperparameters and the need for more efficient gateway selection strategies in federated learning settings with heterogeneous data. The MSEAvg algorithm also increases computational costs due to additional server-side computations. The number of cyber threats is still limited with only Mirai and Gafgyt botnets. Future work should focus on implementing an experimental end-to-end framework to evaluate performance in practical networks, integrating techniques like hierarchy learning to enhance the aggregation algorithm, and optimizing hyperparameters for real-world applications. Furthermore, there should be additional experiments on various IoT network datasets. Addressing these challenges can further refine the proposed framework to better meet the evolving security needs of IoT networks.